OpenStudy (anonymous):

i need help with an integral\[\int x J_n^2(x) \ \text{d}x\]

OpenStudy (anonymous):

@mahmit2012

OpenStudy (anonymous):

\[J _{n}\] is just any arbitary function ?

OpenStudy (anonymous):

bessel function of first kind

OpenStudy (anonymous):

i am out then

OpenStudy (experimentx):

*

OpenStudy (anonymous):

\[x^2\frac{ d^2y }{ dx^2 }+x \frac{ dy }{ dx }+(x^2-n^2)y=0\]

OpenStudy (anonymous):

so the bessel function is the solution to this curve

OpenStudy (anonymous):

can we find the intergral of these

OpenStudy (anonymous):

emm...im just trying !! i dont know :(

OpenStudy (anonymous):

Well my good friend wikipedia tells us that we can expand the Bessel Function as follows with taylor series: http://upload.wikimedia.org/math/1/b/2/1b23400208b273377e8cdec7d82f0242.png

OpenStudy (anonymous):

So then the integral becomes: \[\int x\sum_{m=0}^\infty \frac{(-1)^m}{m!(m+n)!}(\frac{1}{2}x)^{2m+n}\]for integer orders

OpenStudy (anonymous):

\[\int x\left(\sum_{m=0}^\infty \frac{(-1)^m}{m!(m+n)!}(\frac{1}{2}x)^{2m+n}\right)^2 dx\] - forgot the dx

OpenStudy (anonymous):

can u show me the steps with wolfram?

OpenStudy (anonymous):

No, I can't. But you do see the answer, right? My guess is that to get the answer you have to write \((\sum...)^2\) as a single sum using binomial expansion or something, then you can integrate it easily since it's just a polynomial. Then you reconstruct the sum as three separate sums and get the nice answer that alpha got: http://www.wolframalpha.com/input/?i=integrate+x+BesselJ [n%2C+x]^2+dx

OpenStudy (anonymous):